How to Calculate the Molarity of  13% H2SO4 solution?
 

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
Take exactly 1.0L = 1000mL of this solution 
Mass of this solution = 1020g - from density 

Solution is 13% H2SO4 by mass 
Mass of H2SO4 = 13/100*1020 = 132.6g H2SO4 

molarity : 
Molar mass H2SO4 = 98g/mol 
mol H2SO4 in 132.6g = 132.6/98 = 1.35 mol H2SO4 dissolved in 1.0L solution 
Molarity = 1.35M 
Should be written as 1.4M because of 2 significant figures.
 
 
Regards
Arun (askIITians forum expert)